A simple test of PSNFSS2e in old and new encoding

T1 encoding

\rm  Roman  \itshape  Italic  \scshape  Small Caps

\em  Emphatic \slshape  Slanted  \ttfamily    Typewriter
\bfseries Boldface \sffamily Sans Serif
ò \`{o}  õ \~{o} ǒ \v{o}   o̧ \c{o} ó \'{o}
ō \={o}  ő \H{o}   o. \d{o} ô \^{o} ȯ \.{o}
oo \t{oo}   ō \b{o} ö \"{o} ŏ \u{o}
    \dag  § \S       \begin{matrix}c \\ ◯ \end{matrix}  \copyright ı \i
    \ddag ¶ \P  £  \pounds  ȷ \j
œ \oe  Œ \OE  æ  \ae
Æ \AE  å \aa  Å  \AA
ø \o  Ø \O  ł  \l
Ł \L  ß \ss  ¿  ?`
¡ !`  . . . \ldots L^{A}TEX \LaTeX

This costs ‘£’ 45 to do; this costs £45 to do; this costs £45 in sans italic to do; this costs “£” 45 to do; this costs £45 to do this costs £45 to do for a fée. And so on. The sūn has got his hāt in, hip hip hip hooray! ¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS! — ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

OT1 encoding

Note: do not expect the pounds sign to be quite correct — it will always be in italic, and will be a dollar in the sans font.

\rm  Roman  \itshape  Italic  \scshape  Small Caps

\em  Emphatic \slshape  Slanted  \ttfamily    Typewriter
\bfseries Boldface \sffamily Sans Serif
ò \`{o}  õ \~{o} ǒ \v{o}   o̧ \c{o} ó \'{o}
ō \={o}  ő \H{o}   o. \d{o} ô \^{o} ȯ \.{o}
oo \t{oo}   ō \b{o} ö \"{o} ŏ \u{o}
† \dag  § \S       \begin{matrix}c \\ ◯ \end{matrix}  \copyright ı \i
‡ \ddag ¶ \P  $  \pounds  ȷ \j
œ \oe  Œ \OE  æ  \ae
Æ \AE  å \aa  Å  \AA
ø \o  Ø \O  �  \l
� \L  ß \ss  ¿  ?`
¡ !`  . . . \ldots L^{A}TEX \LaTeX

This costs ‘$’ 45 to do (or $45 even); this costs $45 to do; this costs $45 in sans italic to do; this costs “$” 45 to do; this costs $45 to do this costs $45 to do for a fée. And so on. The sūn has got his hāt in, hip hip hip hooray! ¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS! — ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?